# To reset or not to reset in a finite domain: that is the question

### v1

@@ Abstract
$M(x_{T},x)$ $\equiv\left.\frac{\delta\tau_{0}(x_{T})}{\delta r(x)}\right|_{r(x)\equiv 0},$ (9)
instead of the standard derivative in Eq. (4). This can be computed analytically, as detailed in Sec. III of [40]:
@@
where $\Theta$ stands for Heaviside’s step function. The two terms on the right hand side of Eq. (To reset or not to reset in a finite domain: that is the question) have opposite signs: the first is always negative, whereas the second is always positive. Bulk resetting at $x$ is thus beneficial for $x_{T}x<0$ (target and particle positions at opposite sides of the box), whereas it is detrimental for $x_{T}x>0$ and $|x_{T}|>|x|$ (particle between the center and the target). For $x_{T}x>0$ and $|x_{T}|<|x|$, $M(x_{T},x)$ vanishes.
@@
The sign of $\mu(x)$ provides a transparent criterion to elucidate the optimality of the no-resetting strategy in the bulk. It is guaranteed that $\overline{\tau}$ attains a local minimum for $r(x)=0$ if $\mu(x)>0$ for all $x$. Otherwise, the optimal search strategy involves nonvanishing resetting in the bulk. We will now focus on symmetric target distributions, i.e., $P(x_{T})=P(-x_{T})$. Therein, the optimal resetting strategy inherits the symmetry of the problem and it suffices to analyze the symmetrized function $M_{s}(x_{T},x)\equiv[M(x_{T},x)+M(-x_{T},x)]/2$ and its average with respect to $P(x_{T})$ in the interval $[0,1]$.
@@
The symmetrized function $M_{s}(x_{T},x)$, displayed in the bottom panel of Fig. 2, is thus the key quantity to elucidate whether the strategy with no resetting in the bulk is optimal or not. In this sense, it plays a role similar to the function $M(x_{T})$, Eq. (5), shown in the top panel of Fig. 2, for the homogeneous resetting case—but with an additional dependence on the resetting position $x$. For those points $(x_{T},x)$ at which $M_{s}(x_{T},x)<0$, nonresetting in the bulk is not the best strategy for a target perfectly localized at $x_{T}$: introducing an infinitesimal delta kick at $x$ makes $\overline{\tau}$ decrease with respect its value with no resetting. Interestingly, there is an interval of “small” target positions, $0\leq x_{T}\leq x_{T}^{(0)}$, where $M_{s}<0$ for all $x$: this means that the optimal strategy for a target distribution with finite support in $(0,x_{T}^{(0)})$ involves nonvanishing resetting in the bulk regardless of further details of the target distribution. Conversely, there is not an interval of “large” target positions above which $M_{s}$ is positive for all $x$ and, thus, we cannot guarantee that a finite-support distribution leads to suppression of resetting in the bulk as the optimal strategy. However, looking at the figure, it is clear that the target positions that contribute to $r(x)=0$ as the optimal strategy are those close to the boundaries, i.e., $|x_{T}|$ close to unity—as already discussed in the simpler case of homogeneous resetting.
Now, we turn to studying the optimal resetting strategy for the $\beta$-family of target distributions (6). Since it is symmetric, $M_{s}(x_{T},x)$ is the key quantity. We write $\mu(x;\beta)$ because it depends on $\beta$—its analytical expression in terms of hypergeometric functions can be found in Sec. III of [40]. Nevertheless, we focus here on its sign, which is the main property for our purposes. For $\beta<1$, zero bulk resetting is optimal with $\mu(x;\beta)>0$ for $x\in(0,1)$. For $\beta>1$, $\mu(x;\beta)$ becomes negative close to the wall $x=1$, signaling that the no-resetting strategy is no longer the optimal one. Therefore, for the $\beta$-family (6), the flat distribution is the critical case, $\beta_{c}=1$; the critical $\beta$ is consistently smaller compared to the homogeneous case, $\beta_{c}\simeq 1.71$.
@@
The optimal search strategy with bulk resetting cannot be derived analytically for any $P(x_{T})$, because Eq. (3) does not have an analytical solution for arbitrary $r(x)$. We therefore present numerical solutions, based on gradient descent algorithms [41], leading to the optimal $r^{*}(x)$—see Sec. IV of [40]. The obtained optimal strategy displays robust features for $\beta>1$, as shown in Fig. 3: there appears a value $x_{J}$ that separates zero resetting and nonzero resetting regions, i.e., $r^{*}(x)=0$ for $0<|x|<x_{J}$, whereas $r^{*}(x)\neq 0$ for $x_{J}<|x|<1$. Resetting is suppressed in $|x|<x_{J}$ to ensure uninterrupted exploration. For $|x|>x_{J}$, resetting is significantly inhomogeneous, displaying quite high values close to the walls. At $|x|=x_{J}$ the numerical results evidence the presence of a Dirac delta contribution to $r(x)$, which is corroborated by the use of different meshes. Interestingly, we report that the qualitative features of the profile for the optimal resetting obtained, i.e., the vanishing window, the peak and the high heterogeneity, are robust for other families with accumulation of probability in the center, as detailed in Sec. V of [40]. We further find that, for a specific family of target distributions, resetting strategies consisting in two symmetric delta peaks are the exact optimal solution, as detailed in Sec. V of [40].
@@
Figure 3: Optimal heterogeneous resetting profiles in the bulk. Specifically, they have been obtained numerically for the target distributions of the family (6) corresponding to $\beta=0.5,2,3$. For $\beta\leq 1$, the optimal strategy consistently involves no resetting in the bulk. For $\beta>1$, for which nonzero bulk resetting is the optimal strategy, the features displayed—central region with no resetting, strongly heterogeneous behavior close to the wall—are robust. Symbols and lines respectively stand for meshes with $N=501$ and $N=2001$ nodes. The peaks become consistently higher and narrower when increasing the number of nodes in the mesh.
@@
Conclusions.- In this Letter, we have shown that for one-dimensional search problems within a finite domain, resetting boundaries are always beneficial: if the searcher has not found the target upon reaching the wall, it is advantageous to be reset to the initial point. Then, we have analyzed the impact of the resetting strategy in the bulk, taking into account the role of the spatial distribution of the target. We have provided rigorous mathematical criteria to determine the optimality of the absence of resetting in the bulk, for both homogeneous and inhomogeneous resetting rates. These criteria, valid for arbitrary target distributions, have been further illustrated by considering a monoparametric family of such distributions—allowing us to shed further light on the physical reasons behind the optimality/nonoptimality of bulk resetting.
@@
Our work paves new research lines in the study of search processes. On the one hand, we have shown that the spatial distribution of the target plays a key role in determining the optimal search strategy. This contrasts with the typical approach of optimizing the search strategy for a fixed, known position of the target: in an actual search process, the searcher does not know the position of the target. On the other hand, the optimal resetting strategy is highly nontrivial and inhomogeneous, even for the relatively “simple” one-dimensional case analyzed in this Letter. Trying to find an approximate expression for it is a promising research direction, as also is extending our analysis to searches in multidimensional domains.
@@
Acknowledgments.- G. García-Valladares, C. A. Plata and A. Prados acknowledge financial support from Grant ProyExcel_00796 funded by Junta de Andalucía’s PAIDI 2020 programme. C. A. Plata and A. Prados acknowledge financial support from the applied research and innovation Project PPIT2024-31833, cofunded by EU–Ministerio de Hacienda y Función Pública–Fondos Europeos–Junta de Andalucía–Consejería de Universidad, Investigación e Innovación. A. Manacorda acknowledges financial support from the project “MOCA” funded by MUR PRIN2022 Grant No. 2022HNW5YL.

### v2

@@ IV Optimization over heterogeneous resetting
where the functional derivative $M(x_{T},x)$ at fixed target is defined by
$M(x_{T},x)$ $\equiv\left.\frac{\delta\tau_{0}(x_{T})}{\delta r(x)}\right|_{r(x)\equiv 0}$
$=\lim_{\varepsilon\to 0^{+}}\frac{\left.\tau_{0}(x_{T})\right|_{r(x_{0})=\varepsilon\delta(x_{0}-x)}-\left.\tau_{0}(x_{T})\right|_{r(x_{0})=0}}{\varepsilon},$ (15)
which takes the place of the standard derivative $M(x_{T})$ in Eq. (10). The physical meaning of $M(x_{T},x)$ is clear: It measures how the MFPT varies when the nonresetting strategy is locally perturbed at point $x$.
The sign of $\mu(x)$ provides a transparent criterion to elucidate the optimality of the nonresetting strategy in the bulk. On the one hand, the positivity of $\mu(x)$ for all $x$ entails that the average MFPT attains a local minimum for $r(x)=0$, since any perturbation from $r(x)=0$ would lead to a larger average MFPT. On the other hand, the emergence of a subinterval where $\mu(x)$ is negative entails the decrease of the average MFPT for a perturbation such that $r(x)\neq 0$ in that subinterval, implying that $r(x)=0$ does not provide the global minimum of the average MFPT.
The derivative in Eq. (15) can be computed analytically, as detailed in Appendix B:
@@
where $\Theta$ stands for Heaviside’s step function. The two terms on the right hand side of Eq. (IV) have opposite signs: The first is always negative, whereas the second is always positive. Bulk resetting at $x$ is:
  1. 1.
    Beneficial if $x_{T}\,x<0$ (target and particle positions at opposite sides of the box);
  2. 2.
    Detrimental for $x_{T}\,x>0$ and $|x_{T}|>|x|$ (particle between the center and the target);
  3. 3.
    Irrelevant if $x_{T}\,x>0$ and $|x_{T}|<|x|$ and thus $M(x_{T},x)=0$.
In the top panel of Fig. 3, the function $M(x_{T},x)$ in Eq. (IV) is displayed. For $M(x_{T},x)\neq 0$, the sign of $M(x_{T},x)$ is given by the sign of the product $x_{T}\,x$. Hence, some general statements on the stability of the nonresetting strategy in the bulk can be made without further knowledge of the target distribution. Specifically, $r(x)=0$ does not minimize the MFPT for any target distribution with support in the subintervals $x_{T}\in(-1,0)$ or $x_{T}\in(0,1)$. This can be physically understood, because if we know the sign of the target position, the optimal strategy is clearly to forbid exploration of the side of the box opposite to the target’s position, i.e., such that $x\,x_{T}<0$, sending the resetting rate to infinity therein.
@@
We now focus on symmetric target distributions, i.e., $P(x_{T})=P(-x_{T})$. Therein, the optimal resetting strategy inherits the symmetry of the problem and it suffices to analyze the symmetrized function $M_{s}(x_{T},x)\equiv[M(x_{T},x)+M(-x_{T},x)]/2$ and its average with respect to $P(x_{T})$ in the interval $[0,1]$. Taking into account Eq. (IV), we find the symmetrized function
$M_{s}(x_{T},x)=\begin{cases}-\dfrac{1}{4}x^{2}[1+x+(x-3)x_{T}],&0\leq x\leq x_{T},\\ \,\\ -\dfrac{1}{4}(1-x)^{2}x\,x_{T},&x_{T}\leq x\leq 1.\end{cases}$ (17)
The symmetrized function $M_{s}(x_{T},x)$, displayed in the bottom panel of Fig. 3, is thus the key quantity to elucidate whether the strategy with no resetting in the bulk is optimal or not.
@@
Interestingly, there is an interval of “small” target positions, $0\leq x_{T}\leq x_{T}^{(0)}$, where $M_{s}<0$ for all $x$: This means that, for a target distribution with finite support in $(0,x_{T}^{(0)})$, independently of its detailed shape, $\mu$ is negative for all $x$ and the optimal strategy involves nonvanishing resetting in the bulk regardless of further details of the target distribution. Conversely, there is no interval of “large” target positions above which $M_{s}$ is positive for all $x$ and, thus, we cannot guarantee that a finite-support distribution leads, without further knowledge of its details, to suppression of resetting in the bulk as the optimal strategy. However, looking at the figure, it is clear that the target positions that contribute to $r(x)=0$ as the optimal strategy are those close to the boundaries, i.e., $|x_{T}|$ close to unity—as already discussed in the simpler case of homogeneous resetting.
Figure 4: Functional derivative $\mu$ controlling the stability of the nonresetting strategy in the bulk. Specifically, we show $\mu$ as a function of $x$ for the beta family of target distributions introduced in Eq. (12), given in Eq. (18). The critical value $\beta_{c}=1$ (black solid line) marks the change to a not-always-positive $\mu(x;\beta)$, which entails an optimal nonzero resetting strategy in the bulk.
Now, we turn to studying the optimal resetting strategy for the $\beta$-family of target distributions (12). Since it is symmetric, $M_{s}(x_{T},x)$ is the function to be integrated in order to obtain $\mu(x;\beta)$—where we have explicitly introduced the dependence on the $\beta$ parameter. In this case, explicit integration of $M_{s}$ is feasible, with the result
$\mu(x;\beta)=\frac{1}{4}x\Bigg\{$ $-x(1+x)+\frac{\Gamma\left(\frac{1}{2}+\beta\right)}{\sqrt{\pi}\,\Gamma(1+\beta)}$
$\times\Bigg[-(1-x)^{2}+(1-x)^{\beta}(1+x)^{1+\beta}$
$\qquad+2\beta x^{2}(1+x)\,_{2}F_{1}\left(\frac{1}{2},1-\beta;\frac{3}{2},x^{2}\right)\Bigg]\Bigg\}$ (18)
In Fig. 4, we have plotted the functional derivative $\mu$ for different values of $\beta$ as a function of $x$. If for any $x$, $\mu$ takes negative values, this implies that the nonresetting strategy in the bulk is not the optimal one. We have exemplified five values of $\beta$, including the critical value $\beta_{c}=1$ in black, which stands for the flat distribution. For $\beta\leq 1$, zero bulk resetting is optimal with $\mu(x;\beta)>0$ for $x\in(0,1)$. For $\beta>1$, $\mu(x;\beta)$ becomes negative close to the wall $x=1$ (boundary instability), signaling that the nonresetting strategy is no longer the optimal one since $\mu$ becomes negative at least in a certain subinterval. Herein, the critical $\beta$ is consistently smaller compared to the homogeneous case, where we had $\beta_{c}\simeq 1.71$.
V Optimal resetting strategy: Numerical results
The optimal search strategy with bulk resetting cannot be derived analytically for any $P(x_{T})$, because Eq. (7) does not have an analytical solution for arbitrary $r(x)$. We therefore present numerical solutions, based on gradient descent algorithms [49] leading to the optimal $r^{*}(x)$. The numerical optimization is carried out as follows. The spatial coordinate $x$ is discretized into a mesh of $N$ nodes. An initial search strategy $r_{0}(x)$ is considered and the average MFPT is computed for such an initial strategy. The following steps are then executed iteratively:
  1. 1.
    For the current search strategy, the functional derivative $\mu$ is computed in the mesh. Details of the calculation are provided in Appendix C.
  2. 2.
    The new potential search strategy is given by $r(x)\to\max[r(x)-\lambda\,\mu,0]$, where $\lambda$ is an adaptive factor with a given value $\lambda_{0}$ in the first trial.
  3. 3.
    The average MFPT of the new strategy is computed. If it is lower or equal than the previous one, the strategy is updated and we go to step 1. Otherwise, $\lambda$ is reduced, $\lambda\to\lambda/2$, and we go to step 2.
As usual, gradient descent is expected to converge to the local minimum whose attraction basin includes the initial condition, and does not provide any information about global optimality. This issue is related to the nonconvexity of the cost function, which can feature multiple local minima. We mitigate this risk by varying the initial resetting profile, $r_{0}(x)$, either zero or positive uniform values. Numerical parameters typically adopted are $N\in\{501,1001,2001\}$, $\lambda_{0}\in\{10^{5},10^{6},10^{7}\}$. The most consuming task in our gradient descent algorithm is to compute the functional derivative, which is detailed in Appendix C.
@@ V Optimal resetting strategy: Numerical results
In the case of the beta family of target distributions introduced in Eq. (12), the optimal strategy obtained displays robust features as shown in Fig. 5. For $\beta\leq 1$, the numerically optimal resetting strategy converges to vanishing bulk resetting rate as theoretically predicted. For $\beta>1$, there appears a value $x_{J}$ that separates zero resetting and nonzero resetting regions, i.e., $r^{*}(x)=0$ for $0<|x|<x_{J}$, whereas $r^{*}(x)\neq 0$ for $x_{J}<|x|<1$. Resetting is suppressed in $|x|<x_{J}$ to ensure uninterrupted exploration. For $|x|>x_{J}$, resetting is significantly heterogeneous, displaying quite high values close to the walls, i.e., for $|x|\to 1$. At $|x|=x_{J}$ the numerical results evidence the presence of a Dirac delta contribution to $r(x)$, which is corroborated by the use of different meshes—the peak amplitude is proportional to $N$ with a mesh size scaling as $1/N$. Interestingly, we observe that the qualitative features of the profile for the optimal resetting obtained, i.e., the vanishing window, the peak and the high heterogeneity, are robust for other families with accumulation of probability in the center, as detailed in Appendix D.
@@
Figure 5: Optimal heterogeneous resetting profiles in the bulk. Specifically, they have been obtained numerically for the target distributions of the family (12) corresponding to $\beta=0.5,2,3$. For $\beta\leq 1$, the optimal strategy consistently involves no resetting in the bulk. For $\beta>1$, for which nonzero bulk resetting is the optimal strategy, the features displayed—central region with no resetting for $|x|<x_{J}$, strongly heterogeneous behavior close to the wall—are robust. The acceleration factor with respect a vanishing resetting bulk is around 2% and 7% for $\beta=2$ and $\beta=3$ respectively. Symbols and lines respectively stand for meshes with $N=501$ and $N=2001$ nodes.
One of the most remarkable features of our optimal resetting profile is the emergence of Dirac-delta contributions. To the best of our knowledge, this is the first instance of such a resetting strategy. Although peculiar, delta-like resetting for certain points has no physical issues: It can be regarded, for instance, as a limit process in a narrow spatial window, inside which one has a large homogeneous resetting rate. As this spatial window becomes narrower, the large homogeneous resetting rate increases while the area remains constant. A heuristic interpretation of a Dirac-delta resetting and its practical implementation are given in Appendix E. For some target distributions, the optimal resetting strategy consists of two symmetric delta peaks only. Specifically, let us consider the family
$P(x_{T};a)=ax_{T}^{2}+(4-3a)x_{T}^{4}+\frac{7}{30}(-9+8a)x_{T}^{6},$ (19)
with $0<a<(117+\sqrt{8169})/23$. These specific coefficients are chosen to have a monoparametric distribution family, and the interval of $a$ stems from $p(x_{T};a)\geq 0$ for all $0<x_{T}<1$. When the optimal resetting is different from the nonresetting strategy in the bulk, the optimal strategy comprises two symmetric Dirac delta functions at specific positions and with specific intensity that can be numerically derived.
Figure 6: Functional derivative controlling the stability of the nonresetting strategy for the polynomial family in Eq. (19). The critical value $a_{c}$ (black solid line) marks the change where $\mu(x;a)$ is not always positive, indicating an optimal nonzero resetting strategy.
We prove first that the optimal resetting strategy displays a transition for the family of target distributions in Eq. (19) when varying the control parameter $a$. To do that, we have analytically computed $\mu(x;a)$ for this target distribution. The result is a tenth-degree polynomial that is strictly positive in the interval $0<x<1$ for $a<a_{c}\simeq 8.23$ as shown in Fig. 6. In this case, $\mu(x;a)$ for $a>a_{c}$ becomes negative in at least a subinterval of the bulk, $x\in[0,1]$; we call this behavior a bulk instability, as opposed to the boundary instability defining the optimality criteria for the beta distribution.
For $a>a_{c}$, we address a constrained optimization problem for the average MFPT over the family of strategies given by
$r_{\delta}(x)=\frac{u}{x_{R}}[\delta(x-x_{R})+\delta(x+x_{R})].$ (20)
The computation of the average MFPT can be analytically performed, obtaining $\overline{\tau}$ as a function of $x_{R}$ and $u$. Once this result is obtained, numerical optimization can be carried out over $x_{R}$ and $u$. We note that this constrained optimization does not guarantee that Eq. (20) is the optimal strategy, but the analytical computation of the functional derivative for this resetting profile proves that it gives a (local) minimum of the MFPT: The resulting functional derivative is nonnegative, being equal to zero at the boundaries and at $x=\pm x_{R}$, proving that the heterogeneous resetting strategy in Eq. (20) is locally optimal.
Figure 7: Top: Optimal strategy for the polynomial family in Eq. (19). The solid lines correspond to the theoretical optimal strategy, Dirac deltas characterized by $x_{R}^{*}(a)$ and $u^{*}(a)$, whereas the symbols correspond to the numerically obtained optimal solution for different grid meshes. Center and Bottom: Optimal parameters for the resetting strategy. The theoretical parameters (black solid lines) are compared to the numerical parameters obtained by gradient descent optimization using N=201, 501, 2001N=201,\,501,\,2001 (circles, squares and diamonds, respectively).
Furthermore, we have run numerical optimization for several values of $a$, and we have observed the optimal profile to be consistent with our theoretical expectation. For a dense mesh, we obtain a sharply peaked resetting strategy, as shown in the top panel of Fig. 7. In the central and bottom panels, we show that the values of the position of the Dirac-delta and its intensity, measured in the numerical optimization through its area, are consistent—error bars are willingly eluded to improve the visualization but, when taken into account, numerical results are compatible with the theoretical expectations.
VI Conclusions
@@
In this study, we have obtained the optimal resetting strategies for search processes in a confined $1d$ geometry with distributed targets. We have introduced resetting boundaries, which are always beneficial for our goal: If the searcher has not found the target upon reaching the wall, it is advantageous to be reset to the initial point. We have then analyzed the impact of the resetting strategy in the bulk, taking into account the role of the spatial distribution of the target. We have provided rigorous mathematical criteria to determine when the absence of resetting in the bulk is optimal, under both homogeneous and heterogeneous resetting rates. For homogeneous resetting, the convexity of the relevant function guarantees that, if the MFPT has a local minimum for $r=0$, this is indeed the global minimum. For heterogeneous resetting, our approach rigorously determines whether $r(x)=0$ constitutes a local minimum, since no general convexity argument holds in this case. Remarkably, these criteria, valid for arbitrary target distributions, have been further illustrated by considering two monoparametric families of such distributions—allowing us to shed further light on the physical reasons behind the optimality of bulk resetting.
@@
Our work paves new research lines in the study of search processes. On the one hand, we have shown that the spatial distribution of the target plays a key role in determining the optimal search strategy. This contrasts with the typical approach of optimizing the search strategy for a fixed, known position of the target: In an actual search process, the searcher does not know the position of the target. On the other hand, the optimal resetting strategy is highly nontrivial and heterogeneous—including Dirac-delta peaks that have not been previously reported, even for the relatively “simple” one-dimensional case analyzed in this work. Several future research directions arise: (i) to better understand the physical origin of Delta-like contributions in the optimal resetting rate; (ii) to derive analytical expressions for it; (iii) to extend our analysis to distributions with support over the full real axis and (iv) to searches in multidimensional domains.
Acknowledgements.
@@
G. García-Valladares, C. A. Plata and A. Prados acknowledge financial support from Grant PID2024-155268NB-I00 funded by MICIU/AEI/10.13039/501100011033/ FEDER, UE and Grant ProyExcel_00796 funded by Junta de Andalucía’s PAIDI 2020 programme. C. A. Plata and A. Prados acknowledge financial support from the applied research and innovation Project PPIT2024-31833, cofunded by EU–Ministerio de Hacienda y Función Pública–Fondos Europeos–Junta de Andalucía–Consejería de Universidad, Investigación e Innovación. A. Manacorda acknowledges financial support from the project “MOCA” funded by MUR PRIN2022 Grant No. 2022HNW5YL.
Data availability